# Credit Risk Estimation with Non-Financial Features: Evidence from a Synthetic Istanbul Dataset

### v1

@@ 4.1 Effect of alternative data
The contrast between CatBoost Demo and CatBoost Full illustrates the incremental value of the behavioural and asset variables. Adding phone characteristics, subscription spend, online shopping cadence, and car ownership raises AUC by 0.013 and boosts $F_{1}$ by 0.10. A paired DeLong test confirms that the AUC change is significant at $p<0.001$. Similar deltas appear for LightGBM and XGBoost, which each gain about 0.014 AUC and 0.10 $F_{1}$.
These gains echo findings in prior work that mobile phone usage, call detail records, digital footprint metadata, and super app behaviour each add predictive signal beyond basic age, income, and education. Our results extend that evidence by showing that a compact set of nine alternative attributes, readily available from a handset or an e commerce profile, can markedly improve credit risk discrimination for consumers who possess no bureau file.
In operational terms, switching from Demographic only model to Demographic with alternative model would allow a lender to approve roughly eleven additional credit worthy applicants per one hundred screened while rejecting seven more high risk cases, without any reliance on conventional bureau data. These improvements underline the practical value of incorporating alternative signals when assessing the credit worthiness of underbanked populations.
@@ 5 Discussion
Alternative data have already reshaped micro lending in Sub Saharan Africa, where mobile money usage predicts delinquency with surprising accuracy Bjorkegren2020. In Latin America regulators now permit telecom scoring products that combine handset age with call network diversity, a policy move that has unlocked small business funding for street vendors and ride hail drivers. Our Istanbul study confirms the external validity of those earlier findings while adding realism for a large, urban, middle income market where smartphone penetration exceeds ninety percent and e commerce growth reaches double digits annually.
A natural worry is that behavioral proxies might simply replicate socio economic bias. The DICE analysis offers partial reassurance. Still, true fairness cannot be declared until models are stress tested on live portfolios that include migrants, refugees, and workers, groups whose digital patterns may diverge from the synthetic population.
@@
Deep learning remains an open frontier. Text embeddings of phone brand reviews or social media bios might reveal soft skills like conscientiousness or optimism, dimensions shown to correlate with repayment in psychometric studies Arraiz2015Psychometrics. Transformer models fine tuned on telecom sequences could detect subtle churn signals weeks before a payment lapse. These techniques require larger feature canvases and sophisticated privacy guards but promise additional lift for segments where even online shopping data are sparse.
@@ 6 Conclusion
This work set out to answer a practical question: can credit worthiness be estimated with sufficient accuracy when the only information available comes from a person’s phone, e commerce behavior, and a handful of observable traits such as age and residence while ignoring bureau files entirely. Using a synthetic dataset of one hundred thousand Istanbul residents, generated from TÜİK marginals and enriched with domain heuristics, we trained and evaluated twelve model variants that differ in algorithm family and feature scope. The accompanying Jupyter notebook documents every configuration detail, shows how early stopping limits over fitting, and demonstrates how DICE produces sparse yet diverse counter factual explanations.
@@
Empirical results leave little room for doubt. Across CatBoost, LightGBM, and XGBoost the addition of nine alternative variables lifts AUC by about 0.013 and raises balanced $F_{1}$ by ten points. Precision sees the sharpest improvement. CatBoost Full, for example, flags ninety six of one hundred defaulters while issuing fewer than five false approvals in the same batch, a separation level that pure demographic scorecards cannot reach. Decision tree and random forest models also gain, proving that the signal exists even with simpler learners. Logistic Regression shows only minor improvement, underscoring that the new variables interact in nonlinear ways that linear methods cannot capture. Concrete application scenarios highlight the business value. A lender that currently declines every applicant without a bureau file could adopt LightGBM Full, set a threshold that maintains the loss rate of its prime book, and approve roughly fourteen additional loans per hundred underbanked applicants. At Istanbul’s typical micro loan size of fourteen thousand Turkish lira, this represents about two hundred thousand lira in additional credit per hundred screened clients without raising expected loss. Sensitivity analysis in the notebook shows that lowering the cut off by four AUC points still keeps expected loss below three percent, a level acceptable for most instalment products.
@@
The counter factual study adds interpretability. In forty two percent of cases where the prediction flips from reject to accept, changing just one variable online shopping frequency, subscription cost, or phone age is sufficient. This finding echoes field evidence from Latin America and East Africa in which purchase rhythm and handset replacement cycle stand in for disposable income and social capital. It also offers practical guidance: lenders need only a concise set of behavioral indicators, collected with user consent, to gain most of the predictive lift. Validation on ten thousand real, de identified loan files reinforces external relevance. The synthetic trained CatBoost Full model remains calibrated within two points on the live portfolio, and the pattern of DICE explanations is stable, suggesting that the generator preserves key relationships among income, device tier, and spending habits. Although the real sample is smaller and may contain selection bias, this transfer indicates that the synthetic benchmark is a credible proxy for operations data.
@@
Future work should inspect age and income groups and apply causal tools that distinguish genuine repayment capacity from lifestyle signals correlated with protected traits. The alternative block could grow to include bill payment regularity or anonymised mobility entropy, both shown in earlier studies to enhance prediction. Finally, every stage of the pipeline relies on user consent and secure data handling; without strict governance, the benefits of alternative data could turn into privacy risk. Even with these caveats, the study makes three concrete contributions. First, it delivers an open, Istanbul specific synthetic dataset that removes the main barrier researchers face when studying underbanked credit scoring, namely the lack of shareable borrower level data. Second, it provides a fully reproducible modeling pipeline with clear evidence that alternative variables can substitute for bureau files without sacrificing accuracy. Third, it offers regulators and lenders a worked example of how behavioral data can be used responsibly, complete with transparent explanations and documented fairness checks. Taken together, the evidence shows that alternative data, when thoughtfully selected and ethically deployed, can close much of the information gap that excludes millions of credit invisible consumers. Moving beyond legacy bureau variables allows financial institutions to expand access, diversify portfolios, and support inclusive economic growth while maintaining prudent risk management.
Acknowledgements.

### v3

@@ 4.1 Effect of alternative data
The contrast between CatBoost Demo and CatBoost Full illustrates the incremental value of the behavioral and asset variables. Adding phone characteristics, subscription spend, online shopping cadence, and car ownership raises AUC by 0.013 and boosts $F_{1}$ by 0.10. A paired DeLong test confirms that the AUC change is significant at $p<0.001$. Similar deltas appear for LightGBM and XGBoost, which each gain about 0.014 AUC and 0.10 $F_{1}$.
These gains echo findings in prior work that mobile phone usage, call detail records, digital footprint metadata, and super app behavior each add predictive signal beyond basic age, income, and education. Our results extend that evidence by showing that a compact set of ten alternative attributes, readily available from a handset or an e commerce profile, can markedly improve credit risk discrimination for consumers who possess no bureau file.
To translate score improvements into an operational scenario, we evaluate lift at a fixed operating policy. We fix the approval rate at r percent and choose, within each fold, a threshold that approves the top r percent of applicants by predicted score on the training portion of that fold. We then apply that threshold to the held out fold and compute, per 100 screened applicants, the number of additional credit worthy approvals and the number of additional high risk rejections achieved by the Full model relative to the Demo model. We report the mean across folds with a 95 percent bootstrap confidence interval. We also report lift at a fixed default rate target. For each fold, we select a threshold on the training portion that achieves a default rate of t percent among approved applicants, then apply it to the held out fold and compute the same per 100 screened lift quantities.
Confidence intervals use B = 1000 stratified bootstrap resamples drawn at the applicant level from the pooled out of fold predictions, preserving the delinquency_FL prevalence. We report percentile intervals for the mean lift across resamples.
@@ 5 Discussion
Alternative data have already reshaped micro lending in Sub Saharan Africa, where mobile money usage predicts delinquency with surprising accuracy Björkegren and Grissen (2020). In Latin America regulators now permit telecom scoring products that combine handset age with call network diversity, a policy move that has unlocked small business funding for street vendors and ride hail drivers. Our Istanbul study confirms the external validity of those earlier findings while adding realism for a large, urban, middle income market where smartphone penetration exceeds ninety percent and e commerce growth reaches double digits annually. A natural worry is that behavioral proxies might simply replicate socio economic bias. The DICE analysis offers partial reassurance.
@@
Deep learning remains an open frontier. Text embeddings of phone brand reviews or social media bios might reveal soft skills like conscientiousness or optimism, dimensions shown to correlate with repayment in psychometric studies Arráiz et al. (2015). Transformer models fine tuned on telecom sequences could detect subtle churn signals weeks before a payment lapse. These techniques require larger feature canvases and sophisticated privacy guards but promise additional lift for segments where even online shopping data are sparse.
@@ 6 Conclusion
This work set out to answer a practical question: can credit worthiness be estimated with sufficient accuracy when the only information available comes from a person’s phone, e commerce behavior, and a handful of observable traits such as age and residence while ignoring bureau files entirely. Using a synthetic dataset of one hundred thousand Istanbul residents, generated from TÜİK (TURKSTAT) marginals and enriched with domain heuristics, we trained and evaluated twelve model variants that differ in algorithm family and feature scope. The accompanying Jupyter notebook documents every configuration detail, shows how early stopping limits over fitting, and demonstrates how DICE produces sparse yet diverse counter factual explanations.
@@
Empirical results leave little room for doubt. Across CatBoost, LightGBM, and XGBoost the addition of ten alternative variables lifts AUC by about 0.013 and raises balanced $F_{1}$ by ten points. Precision sees the sharpest improvement. CatBoost Full, for example, flags ninety six of one hundred defaulters while issuing fewer than five false approvals in the same batch, a separation level that pure demographic scorecards cannot reach. Decision tree and random forest models also gain, proving that the signal exists even with simpler learners. Logistic Regression shows only minor improvement, underscoring that the new variables interact in nonlinear ways that linear methods cannot capture. Concrete application scenarios highlight the business value. A lender that currently declines every applicant without a bureau file could adopt LightGBM Full, set a threshold that maintains the loss rate of its prime book, and approve roughly fourteen additional loans per hundred underbanked applicants. At Istanbul’s typical micro loan size of fourteen thousand Turkish lira, this represents about two hundred thousand lira in additional credit per hundred screened clients without raising expected loss. Sensitivity analysis in the notebook shows that lowering the cut off by four AUC points still keeps expected loss below three percent, a level acceptable for most installment products.
@@
The counter factual study adds interpretability. In forty two percent of cases where the prediction flips from reject to accept, changing just one variable online shopping frequency, subscription cost, or phone age is sufficient. This finding echoes field evidence from Latin America and East Africa in which purchase rhythm and handset replacement cycle stand in for disposable income and social capital. It also offers practical guidance: lenders need only a concise set of behavioral indicators, collected with user consent, to gain most of the predictive lift. The generator preserves key relationships among income, device tier, and spending habits. Although the real sample is smaller and may contain selection bias, this transfer indicates that the synthetic benchmark is a credible proxy for operations data.
@@
Future work should inspect age and income groups and apply causal tools that distinguish genuine repayment capacity from lifestyle signals correlated with protected traits. The alternative block could grow to include bill payment regularity or anonymized mobility entropy, both shown in earlier studies to enhance prediction. Finally, every stage of the pipeline relies on user consent and secure data handling; without strict governance, the benefits of alternative data could turn into privacy risk. Even with these caveats, the study makes three concrete contributions. First, it delivers an open, Istanbul specific synthetic dataset that removes the main barrier researchers face when studying underbanked credit scoring, namely the lack of shareable borrower level data. Second, it provides a modeling pipeline with clear evidence that alternative variables can substitute for bureau files without sacrificing accuracy. Third, it offers regulators and lenders a worked example of how behavioral data can be used responsibly, complete with transparent explanations. Taken together, the evidence shows that alternative data, when thoughtfully selected and ethically deployed, can close much of the information gap that excludes millions of credit invisible consumers. Moving beyond legacy bureau variables allows financial institutions to expand access, diversify portfolios, and support inclusive economic growth while maintaining prudent risk management.
Acknowledgments